A rectangle is six times as long as it is wide. If the area is 294 ft squared, what is length and width of the table?
Answer by checkley77(12844)

W=6L
LW=AREA
L(6L)=294
6L^2=294
L^2=294/6
L^2=49
L=SQRT49
L=9 ANS.
W=6*7=42 ANS.
PROOF:
42*9=294
294=294